OpenStudy (anonymous):

What are the solutions of the equation? Graph and check the solutions. | x | + 10 = 1

OpenStudy (anonymous):

subtract 10 from both sides; what do you get

OpenStudy (anonymous):

9 :/

OpenStudy (anonymous):

1-10= -9 |x| = -9 can you get a negative number when you take the absolute values?

OpenStudy (anonymous):

+ - - = - ?

OpenStudy (anonymous):

| x | + 10 = 1 subtract 10 | x | + 10 -10 = 1 - 10 | x | = -9 so it will be no solution, because when you have absolute values you always get positive value

OpenStudy (anonymous):

| -x | = | x | = x

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

5 bucks :p